OpenStudy (zuha):

Tanya wants to gift-wrap a container that is in the shape of a right circular cylinder. The radius and height of the container are given below in centimeters. The radius is 2.1 cm and the height is 8.8 cm

OpenStudy (ranga):

The question is missing.

OpenStudy (zuha):

Oops, the question is; which estimate is closest to the minimum amount of gift wrap Tanya needs to cover the entire container?

OpenStudy (ranga):

You need to find the surface area of the cylinder. The lateral surface area is: 2 * pi * r * H The surface area of top and bottom circles are: pi * r^2 + pi * r^2 Plug in the numbers and compute the total surface area.

OpenStudy (zuha):

The estimates r; 120 squared.....132 squared.....432 squared.... 216 squared

OpenStudy (ranga):

Total surface area = 2 * pi * r * H + pi * r^2 + pi * r^2 = 2 * pi * r * H + 2 * pi * r^2 = 2(pi)(r) * (H + r) Plug in the numbers and see what you get.

OpenStudy (ranga):

The number that is greater than what the above calculation yields and closest to the calculation will be your answer.